# DIVER-1: Deep Integration of Vast Electrophysiological Recordings at Scale

### v1

@@ F.2 Finetuning Dataset Description
Neuroprobe Neuroprobe (2024neuroprobe) is a large scale iEEG benchmarks with naturalistic labels during movie watching. 10 subjects watch 25 movies, age from 6 to 19. There are 3 types of evaluation in neuroprobe; single subject-single movie (WithinSession) (splits within the movies), single subject-different movie(CrossSession) (splits within subjects), different subject-different movie(CrossSubject). We evaluated the model in WithinSession. Additionally, Neuroprobe provides an option to subset subjects and trials. We used the LITE option (default configuration), which includes two movies per subject and a total of six subjects.  [^5] PopT’s performance on the same task differs between its original paper and its evaluation in the neuroprobe benchmark because neuroprobe implemented proper train/test splits across time so that no temporal leakage occurs between training and test sets, whereas the original PopT evaluation used random sampling that can lead to data contamination across temporal boundaries. Detailed description of each task is provided below (adapted from (2024neuroprobe):
@@
    FACED (chen2023faced): A large-scale EEG corpus for emotion recognition. It contains recordings from 123 subjects with 32-channel EEG while watching 28 emotion-eliciting video clips. Emotions are categorized into 9 discrete classes: amusement, inspiration, joy, tenderness, anger, fear, disgust, sadness, and neutral. We evaluated the model with the 9-class emotion classification task.
@@
    PhysioNet-MI (goldberger2000physionet; schalk2004bci2000): An EEG dataset for motor imagery–based BCI tasks. It includes recordings from 109 subjects using a 64-channel 10–20 montage and contains four motor imagery classes: left fist, right fist, both fists, and both feet.
@@
    MentalArithmetic (zyma2019electroencephalograms): An EEG dataset for mental stress detection. It contains recordings from 36 subjects using 20 channels. We used 19 channels in total, excluding 1 reference channel. The dataset consists of recordings during mental arithmetic tasks under two conditions: with mental stress and without mental stress.
@@
    Mumtaz2016 (Mumtaz2016): A clinical EEG dataset designed to distinguish major depressive disorder patients from healthy individuals. This dataset comprises 64 subjects (34 with MDD, 30 healthy controls), with signals acquired from 19 scalp locations following the standard 10-20 electrode placement system. We employed the resting-state conditions for binary MDD classification.
@@
    TUEV (obeid2016tueg): An EEG dataset for event type classification in clinical neurophysiology. This corpus provides annotated EEG segments categorized into six classes: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG). We evaluated the model on this 6-class event classification task.
@@ F.3 QAQC and preprocessing
All data underwent quality assessment and control (QAQC) and preprocessing with a philosophy of minimal intervention to retain as much original signal information as possible. For QAQC, we normalized signals by dividing EEG by 100 µV and iEEG by 200 µV (the latter accounting for larger amplitudes in intracranial recordings). While jiang2024labram applied normalization without QAQC and wang2024cbramod removed entire segments if even one timepoint exceeded 100 µV, we adopted a more conservative clipping approach to prevent data loss. We clipped amplitude values exceeding these normalization thresholds, only discarding electrodes when more than 3.33% of samples required clipping and removing whole segments when more than 50% of channels were compromised. This conservative strategy enabled us to preserve substantially more usable data: whereas CBRaMod’s preprocessing yielded approximately 174.7k channel-hours of pretraining data on the same TUEG dataset (refer to Appendix Table 26), our QAQC pipeline retained 422k channel-hours, a 2.4× increase.
@@ Appendix G Comparison With Existing EEG/iEEG Foundation Models
A direct comparison of training epochs for prior EEG/iEEG foundation models is misleading due to varying dataset sizes (Table 25). To enable a fair assessment, we introduce the “Scaled Epochs on Our Data” metric. This normalizes the total data processed during training (channel-hours $\times$ epochs) into an equivalent number of epochs on our dataset, allowing for a direct comparison of training epochs across all prior models and our own. For entries in Table 25 that use estimated values (marked with *), the corresponding estimation procedures are documented in Table 26.
@@
BENDR (kostas2021bendr) EEG 155M* N/A 1 N/A
BrainBERT (wang2023brainbert) SEEG 43M* 4.5k 39* 0.5
Brant (zhang2023brant) SEEG 505.68M 281k 32* 25.6
BIOT (yang2023biot) EEG 3.3M 312k 100 88.6
Neuro–GPT (cui2024neurogpt) EEG 90M* 541k 135 207.6
LaBraM (jiang2024labram) EEG 5.8M, 46M, 369M 76.8–83.7k 50 11.4
EEGPT (wang2024eegpt) EEG 0.4M, 0.5M, 1.6M, 6.4M, 19M, 25M, 76M, 101M 11.1k* 200 6.8
CBraMod (wang2024cbramod) EEG 0.1M, 0.4M, 0.8M, 1.2M, 1.5M, 2M, 3M, 4M 175.7k* 40 20

### v2

@@ F.2 Finetuning Dataset Description
Neuroprobe Neuroprobe (Zahorodnii et al., 2025) is a large scale iEEG benchmarks with naturalistic labels during movie watching. 10 subjects watch 25 movies, age from 6 to 19. There are 3 types of evaluation in neuroprobe; single subject-single movie (WithinSession) (splits within the movies), single subject-different movie(CrossSession) (splits within subjects), different subject-different movie(CrossSubject). We evaluated the model in WithinSession. Additionally, Neuroprobe provides an option to subset subjects and trials. We used the LITE option (default configuration), which includes two movies per subject and a total of six subjects.  [^4] PopT’s performance on the same task differs between its original paper and its evaluation in the neuroprobe benchmark because neuroprobe implemented proper train/test splits across time so that no temporal leakage occurs between training and test sets, whereas the original PopT evaluation used random sampling that can lead to data contamination across temporal boundaries. Detailed description of each task is provided below (adapted from (Zahorodnii et al., 2025):
@@
    FACED (Chen et al., 2023): A large-scale EEG corpus for emotion recognition. It contains recordings from 123 subjects with 32-channel EEG while watching 28 emotion-eliciting video clips. Emotions are categorized into 9 discrete classes: amusement, inspiration, joy, tenderness, anger, fear, disgust, sadness, and neutral. We evaluated the model with the 9-class emotion classification task.
@@
    PhysioNet-MI (Goldberger et al., 2000; Schalk et al., 2004): An EEG dataset for motor imagery–based BCI tasks. It includes recordings from 109 subjects using a 64-channel 10–20 montage and contains four motor imagery classes: left fist, right fist, both fists, and both feet.
@@
    MentalArithmetic (Zyma et al., 2019): An EEG dataset for mental stress detection. It contains recordings from 36 subjects using 20 channels. We used 19 channels in total, excluding 1 reference channel. The dataset consists of recordings during mental arithmetic tasks under two conditions: with mental stress and without mental stress.
@@
    Mumtaz2016 (Mumtaz, 2016): A clinical EEG dataset designed to distinguish major depressive disorder patients from healthy individuals. This dataset comprises 64 subjects (34 with MDD, 30 healthy controls), with signals acquired from 19 scalp locations following the standard 10-20 electrode placement system. We employed the resting-state conditions for binary MDD classification.
@@
    TUEV (Obeid and Picone, 2016): An EEG dataset for event type classification in clinical neurophysiology. This corpus provides annotated EEG segments categorized into six classes: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG). We evaluated the model on this 6-class event classification task.
@@ F.3 QAQC and preprocessing
[…]
@@ Appendix G Comparison With Existing EEG/iEEG Foundation Models
A direct comparison of training epochs for prior EEG/iEEG foundation models is misleading due to varying dataset sizes (Table  27). To enable a fair assessment, we introduce the “Scaled Epochs on Our Data” metric. This normalizes the total data processed during training (channel-hours $\times$ epochs) into an equivalent number of epochs on our dataset, allowing for a direct comparison of training epochs across all prior models and our own. For entries in Table 27 that use estimated values (marked with *), the corresponding estimation procedures are documented in Table 28.
@@
BENDR (Kostas et al., 2021) EEG 155M* N/A 1 N/A
BrainBERT (Wang et al., 2023) SEEG 43M* 4.5k 39* 0.5
Brant (Zhang et al., 2023) SEEG 505.68M 281k 32* 25.6
BIOT (Yang et al., 2023) EEG 3.3M 312k 100 88.6
Neuro–GPT (Cui et al., 2024) EEG 90M* 541k 135 207.6
LaBraM (Jiang et al., 2024) EEG 5.8M, 46M, 369M 76.8–83.7k 50 11.4
EEGPT (Wang et al., 2024b) EEG 0.4M, 0.5M, 1.6M, 6.4M, 19M, 25M, 76M, 101M 11.1k* 200 6.8
CBraMod (Wang et al., 2024c) EEG 0.1M, 0.4M, 0.8M, 1.2M, 1.5M, 2M, 3M, 4M 175.7k* 40 20